# A highly efficient single-loop smoothing damped Newton method for large-scale bilevel hyperparameter optimization of SVC

### v1

@@ Abstract
Bilevel hyperparameter optimization has received growing attention thanks to the fast development of machine learning. Due to the tremendous size of data sets, the scale of bilevel hyperparameter optimization problem could be extremely large, posing great challenges in designing efficient numerical algorithms. In this paper, we focus on solving the large-scale mathematical programs with equilibrium constraints (MPEC) derived from hyperparameter selection of L1-support vector classification (L1-SVC). We propose a highly efficient single-loop smoothing damped Newton method (SDNM) for solving such MPEC. Compared with most existing algorithms where subproblems are involved and solved by on-shelf packages, our approach fully takes advantage of the structure of MPEC and therefore is single-loop. Moreover, the proposed SDNM enjoys a quadratic convergence rate under proper assumptions. Extensive numerical results over LIBSVM dataset show the superior performance of SDNM over other state-of-art algorithms including the Scholtes global relaxation method (SGRM) with subproblem solved by SNOPT and the Matlab built-in function fmincon, especially in CPU time. For example, for dataset w4a, SDNM is 20 times faster than SGRM and 3 times faster than fmincon. Further numerical results also verifies the quadratic convergence rate of SDNM as well as the fulfillment of the second order sufficient condition, while guarantees that SDNM returns a strict local minimizer of the smoothing problem of MPEC.
@@ 1 Introduction
Bilevel hyperparameter optimization has received growing attention thanks to the fast development of machine learning. Due to the tremendous size of data sets, the scale of bilevel hyperparameter optimization problem could be extremely large, posing great challenges in designing efficient numerical algorithms. Having been one of the most active research areas, vast literature concerns bilevel hyperparameter optimization (BHO) in machine learning. We refer to [1, 8, 9, 16, 30] for surveys and monographs. Below we focus on numerical algorithms for BHO in support vector machine (SVM), which is most relevant to our research in this paper.
@@
We start with an early work by Mangasarian [22], which provides a foundation for an equivalent reformulation (linear-programming problem with complementarity constraints, LPCC) of the so-called counting loss $\|(\cdot)_{+}\|_{0}$. The story begins with Bennett et al.’s work [2], where they propose a bilevel optimization framework for BHO in SVM. By replacing the lower-level problem by KKT conditions to get a single-level problem (which is a mathematical program with equilibrium constraints, MPEC), they proposed a successive linearization method to solve such MPEC, with subproblem solved by CPLEX. Meanwhile, with the equivalent LPCC reformulation of counting loss function, Kunapuli et al. [15] studied BHO for $l_{1}$-regularized support vector classification ($L_{1}$-SVC), with counting loss and $l_{1}$ loss as upper-level functions, respectively. They use FILTER package to solve the resulting MPEC. Later work includes rectangle search algorithm by Lee et al. [17] to solve the resulting single-level LPCC with subproblem also solved by CPLEX, stochastic gradient method by Couellan and Wang [7] for the resulting single-level MPEC (no subproblem involved), and Scholtes global relaxation method [18, 20, 5] also for MPEC, with subproblem solved by SNOPT [12].
@@
Different from above existing work, Wang and Li [27] studied BHO for Logistic-loss SVC with Logistic loss as upper-level and lower-level loss function. Different from BHO for $L_{1}$-SVC and $L_{2}$-SVC, the resulting single-level problem is a nonlinear programming problem (NLP) with only equality constraints. A smoothing Newton method is proposed to solve the KKT conditions of such NLP, which does not involve subproblem. Very recently, Ye et al. [29] propose a difference-of-convex algorithm for BHO in L1-SVM with $l_{1}$-loss upper-level function, with subproblem solved by SDPT3. Similar idea is also studied in [10, 11]. Qian et al. [19] propose an $L_{p}$-Newton method to solve the single MPEC in [15], with subproblem solved by Gurobi.
@@
To summarize, we obtain the following observations. (i) Traditional single-level problem derived from BHO for SVM is based on replacing lower problem by KKT conditions, and leads to an MPEC. (ii) Most algorithms for solving such MPEC are two-loop, meaning that subproblems are involved. (iii) Subproblems are often solved by software packages. However, calling software packages has several drawbacks. Firstly, it ignores the structure of the subproblem, therefore, may not be as efficient as specially designed algorithms for subproblems and may take longer time in solving subproblems. Secondly, with the two-loop setting, the inefficient solving of subproblem will further slow down the progress of outer loop, which lead to embarrassingly longer time for the whole algorithm. Finally, recalling the large-scale of MPEC due to the huge size of data sets, the speed of the algorithms for BHO in SVM is still far from satisfactory. For example, based on our computational experience in [27], the CPU time for the Scholtes global relaxation method (SGRM) in [18] is over one hour when the number of features of the dataset is over 100.
@@
Based on the above analysis, together with our successful experience in designing smoothing Newton method to solve BHO for Logistic-loss SVC, a natural question is whether it is possible to design a specialized single-loop algorithm for MPEC derived from BHO for SVC. This motivate the work in this paper. We focus on designing a fast algorithm for solving single-level MPEC derived from BHO for $L_{1}$-SVC. The contribution of the paper is summarized in three folds. Firstly, we smooth the complementarity constraints in MPEC, and present the perturbation analysis to the resulting nonlinear programming problem with equilibrium constraints (NLPEC). Secondly, we analyze the constraint qualification and the first and second order optimality condition of NLPEC. Next, we design an efficient damped Newton method to solve the KKT condition of NLPEC and show its global convergence and quadratic convergence rate. Finally, we conduct extensive numerical experiments to verify the high efficiency of the proposed smoothing damped Newton method, especially in the superior performance of speed.
Here we would like to highlight the advantage of the proposed SDNM. Firstly, it is single-loop, and does not involve solving subproblem. Secondly, it is software independent, which do not rely on other packages. Finally, by smoothing the complementarity constraints, the number of constraints is reduced to one third, which is much smaller than that solved by other state-of-art methods, and also contributes to the speed up of our algorithm.
@@
The paper is organized as follows. In Sect. 2, we briefly restate the BHO for SVC and the resulting single-level MPEC in [18]. In Sect. 3, we introduce smoothing technique to deal with complementarity constraints in MPEC resulting in a nonlinear programming problem with equilibrium constraints (NLPEC), and analyze the perturbation property. In Sect. 4, we study the theoretical property of NLPEC, including the linear independence constraint qualification (LICQ), the first and second order optimality conditions. In Sect. 5, we present the damped Newton method to solve the KKT conditions of NLPEC, which are essentially nonlinear equations. We also present the global convergence and quadratic convergence rate under reasonable assumptions. In Sect. 6, we conduct extensive numerical experiments to verify the high efficiency of the proposed smoothing damped Newton method (SDNM), especially in the superior performance of speed. Finally, we conclude the paper in Sect. 7.
Notations. For $x\in\mathbb{R}^{n}$, $\|x\|_{0}$ denotes the number of nonzero elements in $x$, while $\|x\|_{2}$ corresponds to $l_{2}$-norm of $x$. $[k]$ denotes $\{1,2,\cdots,k\}$ for the positive integer $k$. |$\arrowvert\Omega\arrowvert$ | denotes the number of elements in the set $\Omega$. We use ‘:=’ to represent ‘define’. We use $\mathbf{1}$ to denote a vector where all elements are ones, and $\mathcal{I}$ to denote the identity matrix. We use Diag$(a_{1},\cdots,a_{n})$ to denote the diagonal matrix with diagonal elements $a_{1},\cdots,a_{n}$. Similarly, we use Diag$(A_{1},\cdots,A_{n})$ to denote a block diagonal matrix with $A_{1},\cdots,A_{n}$ as diagonal block matrices.
2 Bilevel hyperparameter optimization for SVC
In this section,we briefly state BHO for $L_{1}$-SVC and its single-level MPEC, which is the problem that we are going to solve in this paper. More details of the bilevel model and the derivation can be found in [18].
@@ 2 Bilevel hyperparameter optimization for SVC
The bilevel model for hyperparameter selection in $L_{1}$-SVC is based on cross-validation (CV). Consider a dataset $\Omega\in\left\{\left(x_{i},y_{i}\right)\right\}_{i=1}^{l_{1}}\subseteq\mathbb{R}^{n+1}$, where $x_{i}\in\mathbb{R}^{n}$ denotes a data point and $y_{i}\in\left\{\pm 1\right\}$ is the corresponding label. CV involves partitioning the dataset $\Omega$ into $T$ folds to evaluate the performance of $L_{1}$-SVC by dividing the data set into $T$ folds of equal size, denote as $\Omega_{t},\ t\in[T]$. In the $t$-th lower-level problem, we train the $t$-th fold training set $\overline{\Omega}_{t}:=\Omega\backslash\Omega_{t}$ by the soft-margin SVC model with the $l_{1}$-loss function [^1] We choose the $L_{1}$-SVC model as the typical lower-level problem due to the following reasons. Firstly, from the practical perspective, the $L_{1}$-SVC model is a widely used statistical model in machine learning [28, 25]. Secondly, the $L_{1}$-SVC model is more challenging to tackle than the $L_{2}$-SVC model due to the nonsmoothness of the $l_{1}$-loss function.. That is, given $C\geq 0$, we solve the $L_{1}$-SVC model.
The aim of the upper-level problem is to minimize the $T$-fold cross-validation error (CV error) measured on the validation sets based on the optimal solutions of the lower-level problems. Specifically, the basic bilevel optimization model for selecting the hyperparameter $C$ in SVC is formulated as
@@
$\begin{array}[]{cl}\min\limits_{C\in\mathbb{R},\ w^{t}\in\mathbb{R}^{n},\ t=1,\cdots,T}&\frac{1}{T}\sum_{t=1}^{T}\frac{1}{m_{1}}\sum_{i\in\mathcal{N}_{t}}\|\left(-y_{i}\left(x_{i}^{\top}w^{t}\right)\right)_{+}\|_{0}\\ \hbox{s.t.}&C\geq 0,\\ &\text{and for}\ t=1,\cdots,T:\\ &\qquad w^{t}\in\underset{w\in\mathbb{R}^{n}}{\operatorname{argmin}}\left\{\frac{1}{2}\|w\|_{2}^{2}+C\sum_{i\in\overline{\mathcal{N}}_{t}}\left(1-y_{i}\left(x_{i}^{\top}w\right)\right)_{+}\right\},\end{array}$ (1)
@@
where $\mathcal{N}_{t}$ and $\overline{\mathcal{N}_{t}}$ denote the index set of $\Omega_{t}$ and $\overline{\Omega_{t}}$ respectively, and $m_{1}=\arrowvert\Omega_{t}\arrowvert,m_{2}=\arrowvert\overline{\Omega_{t}}\arrowvert$ |. Here, the expression $\sum_{i\in\mathcal{N}_{t}}\|\left(-y_{i}\left(x_{i}^{\top}w^{t}\right)\right)_{+}\|_{0}$ basically counts the number of data points that are misclassified in the validation set $\Omega_{t}$, while the outer summation (i.e., the objective function in (1)) averages the misclassification error over all the folds.
@@
Note that each lower-level problem is strongly convex, one can replace the lower-level problem by its KKT conditions after introducing slack variable $\zeta^{t}\in\mathbb{R}^{m_{2}},\ t=1,\cdots,T$. For the upper level objective function in (1), one can introduce the technique in [22] to rephrase it as a linear programming problem with complementarity constraints [^2] To make the problem well-defined, we assume that $x_{i}^{\top}w^{t}\neq 0$, for all $i\in\mathcal{N}_{t}$ amd $t\in[T]$.. The above two strategies lead to the following MPEC reformulation of (1) (See [18] for more details.),

### v3

# An Efficient Smoothing Damped Newton Method for Large-Scale Mathematical Programs with Equilibrium Constraints

@@ Abstract
Bilevel hyperparameter optimization has received growing attention thanks to the fast development of machine learning. Due to the tremendous size of data sets, the scale of bilevel hyperparameter optimization problem could be extremely large, posing great challenges in designing efficient numerical algorithms. In this paper, we focus on solving the large-scale mathematical programs with equilibrium constraints (MPEC) derived from hyperparameter selection of L1 support vector classification (L1-SVC). We propose a highly efficient smoothing damped Newton method (SDNM) for solving such MPEC. Compared with most existing algorithms where subproblems are solved by packages, our approach fully takes advantage of the structure of MPEC and therefore is package-free. Moreover, the proposed SDNM converges to C-stationary point under MPEC-LICQ with subproblem enjoys a quadratic convergence rate under proper assumptions. Extensive numerical results over LIBSVM dataset show the superior performance of SDNM over other state-of-art algorithms.
@@ 1 Introduction
Bilevel hyperparameter optimization has received growing attention thanks to the fast development of machine learning. Due to the tremendous size of data sets, the scale of bilevel hyperparameter optimization problem could be extremely large, posing great challenges in designing efficient numerical algorithms. Having been one of the most active research areas, vast literature concerns bilevel hyperparameter optimization (BHO) in machine learning. We refer to [9, 35, 46, 48] for surveys and monographs. Below we focus on numerical algorithms for BHO in support vector machine (SVM), which is most relevant to our research in this paper.
@@
We start with an early work by Mangasarian [31], which provides a foundation for an equivalent reformulation (linear-programming problem with complementarity constraints, LPCC) of the so-called counting loss $\|(\cdot)_{+}\|_{0}$. The story begins with Bennett et al.’s work [1], where they proposed a bilevel optimization framework for hyperparameter selection problem in SVM. By replacing the lower-level problem by KKT conditions to get a single-level problem (which is a mathematical program with equilibrium constraints, MPEC), Bennett et al. proposed a successive linearization method to solve such MPEC, with subproblem solved by CPLEX. Meanwhile, with the equivalent LPCC reformulation of counting loss function, Kunapuli et al. [20] studied BHO for $L_{1}$-regularized support vector classification ($L_{1}$-SVC), with counting loss and $L_{1}$ loss as upper-level and lower-level functions, respectively. They used FILTER package to solve the resulting MPEC. Later work includes rectangle search algorithm by Lee et al. [21] to solve the resulting single-level LPCC with subproblem solved by CPLEX, stochastic gradient method by Couellan and Wang [7] for the resulting single-level MPEC (no subproblem involved), and Scholtes global relaxation method [6, 23, 27] for MPEC with subproblem solved by SNOPT [15].
@@
Different from above existing work, Wang and Li [43] studied BHO for Logistic-loss SVC with Logistic loss as upper-level and lower-level loss function. Different from BHO for $L_{1}$-SVC and $L_{2}$-SVC, the resulting single-level problem is a nonlinear programming problem (NLP) with only equality constraints. A smoothing Newton method is proposed to solve the KKT conditions of such NLP, which does not involve subproblem. Very recently, Ye et al. [45] proposed a difference-of-convex algorithm for BHO in $L_{1}$-SVM with $L_{1}$-loss upper-level function, with subproblem solved by SDPT3. Similar idea is also studied in [12, 13]. Qian et al. [24] proposed an $L_{p}$-Newton method to solve the single MPEC in [20], with subproblem solved by Gurobi.
@@
To summarize, we obtain the following observations. (i) Traditional single-level problem derived from BHO for SVM is based on replacing lower problem by KKT conditions, and leads to an MPEC. (ii) Most algorithms for solving such MPEC are two-loop, where subproblems are often solved by software packages. However, calling software packages has several drawbacks. Firstly, it ignores the structure of the subproblem, therefore, may not be as efficient as specially designed algorithms for subproblems and may take longer time in solving subproblems. Secondly, recalling the large-scale of MPEC due to the huge size of data sets, the speed of the algorithms for BHO in SVM is still far from satisfactory. For example, based on our computational experience in [43], the CPU time for the Scholtes global relaxation method (SGRM) in [23] is over one hour when the number of features of the dataset is over 100.
@@
Based on the above analysis, together with our successful experience in designing smoothing Newton method to solve NLP derived from BHO for Logistic-loss SVC, a natural question is whether it is possible to design a specialized algorithm to solve MPEC derived from BHO for $L_{1}$-SVC. This motivates the work in this paper. The contribution of the paper is summarized in three folds. Firstly, we propose a smoothing method to solve large-scale MPEC derived from BHO for $L_{1}$-SVC and the global convergence is established under reasonable assumptions. Secondly, to solve the subproblem, we design the damped Newton method, and analyze the first and second order optimality conditions. Under certain assumptions, we show the global convergence and local quadratic convergence rate. Finally, we conduct extensive numerical experiments to verify the efficiency of the proposed smoothing damped Newton method, especially the superior performance in speed.
@@
The paper is organized as follows. In Section 2, we briefly restate the BHO for $L_{1}$-SVC and the resulting single-level MPEC in [23], which is the problem that we will solve in this paper. In Section 3, we introduce the smoothing approach to deal with complementarity constraints in MPEC and analyze the convergence of the smoothing method. In Section 4, we analyze the perturbation property of the feasible set of MPEC. In Section 5, we propose the damped Newton method to solve the subproblem, and establish its convergence results under certain assumptions. In Section 6, we conduct extensive numerical experiments to verify the high efficiency of the proposed smoothing damped Newton method (SDNM), especially the superior performance in speed. Finally, we conclude the paper in Section 7.
[…]
2 Large-Scale MPEC: A Motivating Example
In this paper, we focus on the following MPEC:
$\begin{array}[]{cl}\min\limits_{v\in\mathbb{R}^{m+1}}&f(v)\\ \hbox{s.t.}&G(v)\geq 0,\ H(v)\geq 0,\ G(v)^{\top}H(v)=0.\end{array}$ (MPEC)
where $f:\mathbb{R}^{m+1}\to\mathbb{R}$, $G:\mathbb{R}^{m+1}\to\mathbb{R}^{m}$, $H:\mathbb{R}^{m+1}\to\mathbb{R}^{m}$ are affine functions. Such form of (MPEC) is motivated by the single-level reformulation of BHO for $L_{1}$-SVC, which we will briefly introduce below. More details can be found in [23].
@@ 2 Large-Scale MPEC: A Motivating Example
The bilevel model for hyperparameter selection in $L_{1}$-SVC is based on cross-validation (CV). Consider a dataset $\Omega\in\left\{\left(x_{i},y_{i}\right)\right\}_{i=1}^{p_{1}}\subseteq\mathbb{R}^{n+1}$, where $x_{i}\in\mathbb{R}^{n}$ denotes a data point and $y_{i}\in\left\{\pm 1\right\}$ is the corresponding label. CV evaluates the performance of $L_{1}$-SVC by dividing the data set into $T$ folds of equal size, denote as $\Omega_{t},\ t\in[T]$. In the $t$-th lower-level problem, given $C\geq 0$, we train the $t$-th fold training set $\overline{\Omega}_{t}:=\Omega\backslash\Omega_{t}$ by the soft-margin SVC model with the $L_{1}$-loss function [^1] We choose the $L_{1}$-SVC model as the typical lower-level problem due to the following reasons. Firstly, from the practical perspective, the $L_{1}$-SVC model is a widely used statistical model in machine learning [8, 44, 49]. Secondly, the $L_{1}$-SVC model is more challenging to tackle than the $L_{2}$-SVC model due to the nonsmoothness of the $L_{1}$-loss function.. The aim of the upper-level problem is to minimize the $T$-fold cross-validation error (CV error) measured on the validation sets based on the optimal solutions of the lower-level problems. Specifically, the basic bilevel optimization model for selecting the hyperparameter $C$ in $L_{1}$-SVC is formulated as
@@
$\small\begin{array}[]{cl}\min\limits_{C\in\mathbb{R},\ w^{t}\in\mathbb{R}^{n},\ t=1,\cdots,T}&\frac{1}{T}\sum\limits_{t=1}^{T}\frac{1}{m_{1}}\sum\limits_{i\in\mathcal{N}_{t}}\|\max\left(0,-y_{i}\left(x_{i}^{\top}w^{t}\right)\right)\|_{0}\\ \hbox{s.t.}&C\geq 0,\\ &\text{and for}\ t=1,\cdots,T:\\ &\hskip 18.49988ptw^{t}\in\underset{w\in\mathbb{R}^{n}}{\operatorname{argmin}}\left\{\frac{1}{2}\|w\|_{2}^{2}+C\sum\limits_{i\in\overline{\mathcal{N}_{t}}}\max\left(0,1-y_{i}\left(x_{i}^{\top}w\right)\right)\right\},\end{array}$ (1)
@@
where $\mathcal{N}_{t}$ and $\overline{\mathcal{N}_{t}}$ denote the index set of $\Omega_{t}$ and $\overline{\Omega_{t}}$ respectively, and $m_{1}$ and $m_{2}$ are the numbers of data points in $\Omega_{t}$ and $\overline{\Omega_{t}}$. Here, the expression $\|\max\left(0,-y_{i}\left(x_{i}^{\top}w^{t}\right)\right)\|_{0}$ basically counts the number of data points that are misclassified in the validation set $\Omega_{t}$, while the outer summation (i.e., the objective function in (1)) averages the misclassification error over all the folds.
@@
Note that each lower-level problem is strongly convex, one can replace the lower-level problem by its KKT conditions after introducing slack variable $\zeta^{t}\in\mathbb{R}^{m_{2}},\ t=1,\cdots,T$. For the upper level objective function in (1), one can introduce the technique in [31] to rephrase it as a linear programming problem with complementarity constraints [^2] To make the problem well-defined, we assume that $x_{i}^{\top}w^{t}\neq 0$, for all $i\in\mathcal{N}_{t}$ amd $t\in[T]$.. The above two strategies lead to the following MPEC reformulation of (1) (see [23] for more details),
@@
where 1 is the vector with all elements one, and
$\small\begin{array}[]{c}A:={\rm Diag}\left(A^{1},\ A^{2},\cdots,A^{n}\right)\in\mathbb{R}^{Tm_{1}\times Tn},\ B:={\rm Diag}\left(B^{1},\ B^{2},\cdots,B^{n}\right)\in\mathbb{R}^{Tm_{1}\times Tn},\\ A^{t}:=\left[\begin{array}[]{c}y_{t_{1}}x_{t_{1}}^{\top}\\ \vdots\\ y_{t_{m_{1}}}x_{t_{m_{1}}}^{\top}\end{array}\right]\in\mathbb{R}^{m_{1}\times n},\ B^{t}:=\begin{bmatrix}y_{t_{m_{1}+1}}x_{t_{m_{1}+1}}^{\top}\\ \vdots\\ y_{t_{p_{1}}}x_{t_{p_{1}}}^{\top}\end{bmatrix}\in\mathbb{R}^{m_{2}\times n},\ t\in[T],\\ {{\zeta:=\left[\begin{array}[]{l}\zeta^{1}\\ \zeta^{2}\\ \vdots\\ \zeta^{T}\end{array}\right]\in\mathbb{R}^{Tm_{1}},\ z:=\left[\begin{array}[]{l}z^{1}\\ z^{2}\\ \vdots\\ z^{T}\end{array}\right]\in\mathbb{R}^{Tm_{1}},\ \alpha:=\left[\begin{array}[]{l}\alpha^{1}\\ \alpha^{2}\\ \vdots\\ \alpha^{T}\end{array}\right]\in\mathbb{R}^{Tm_{2}},\ \xi:=\left[\begin{array}[]{l}\xi^{1}\\ \xi^{2}\\ \vdots\\ \xi^{T}\end{array}\right]\in\mathbb{R}^{Tm_{2}}.}}\end{array}$
Therefore, (2) can be written into a compact form of (MPEC), where $m=2T(m_{1}+m_{2}),\ f(v):=\frac{1}{Tm_{1}}\mathbf{1}^{\top}\zeta$, and $G(v)\in\mathbb{R}^{m},\ H(v)\in\mathbb{R}^{m},\ v\in\mathbb{R}^{m+1}$ are defined by
$G(v):=\left[\begin{array}[]{c}\zeta\\ z\\ \alpha\\ \xi\end{array}\right],\ H(v):=\left[\begin{array}[]{c}AB^{\top}\alpha+z\\ \mathbf{1}-\zeta\\ BB^{\top}\alpha-\mathbf{1}+\xi\\ C\mathbf{1}-\alpha\end{array}\right],\ v:=\left[\begin{array}[]{c}C\\ \zeta\\ z\\ \alpha\\ \xi\end{array}\right].$ (3)
Specifically, one can notice that the number of components in $v$ and the number of complementarity constraints are $m+1$ and $m$, respectively, where $m=2T(m_{1}+m_{2})$, $m_{1}$ and $m_{2}$ are the number of training data and validation data, respectively. When the number of cross validation or the number of dataset increases, it will lead to the increase of the scale of (MPEC). Solving such a large-scale bilevel hyperparameter selection problem for SVC is extremely challenging, in which the cross-validation process increases the scale of the problem exponentially.
We close this section by the following definitions of constraint qualification and stationary point of (MPEC). Let $v^{*}$ be a feasible point for (MPEC), define the following index sets:
$\begin{array}[]{ll}I_{0+}(v^{*})&=\{i\in[m]:G_{i}(v^{*})=0,\,H_{i}(v^{*})>0\},\\ I_{+0}(v^{*})&=\{i\in[m]:G_{i}(v^{*})>0,\,H_{i}(v^{*})=0\},\\ I_{00}(v^{*})&=\{i\in[m]:G_{i}(v^{*})=0,\,H_{i}(v^{*})=0\}.\end{array}$
We usually use $I_{0+}$ instead of $I_{0+}(v^{*})$ when the dependence of $I_{0+}$ on $v^{*}$ is very clear. It is the same with $I_{+0}$ and $I_{00}$.
Definition 2.1.
[jiani30, page 920] (MPEC-LICQ) A feasible point $v^{*}$ for problem (MPEC) satisfies the MPEC-LICQ if and only if the set of gradient vectors
$\left\{\nabla G_{i}(v^{*})\mid i\in I_{0+}\cup I_{00}\right\}\cup\left\{\nabla H_{i}(v^{*})\mid i\in I_{+0}\cup I_{00}\right\}$
is linearly independent.
Definition 2.2.
Let $v^{*}$ be feasible for (MPEC). Then $v^{*}$ is said to be
(a) weakly stationary, if there are multipliers $\gamma$, $\nu\in\mathbb{R}^{m}$ such that
$\nabla f(v^{*})-\sum_{i=1}^{m}\gamma_{i}\nabla G_{i}(v^{*})-\sum_{i=1}^{m}\nu_{i}\nabla H_{i}(v^{*})=0$
and $\gamma_{i}=0\ (i\in I_{+0})$, $\nu_{i}=0\ (i\in I_{0+})$;
(b) C-stationary, if it is weakly stationary and $\gamma_{i}\nu_{i}\geq 0$ for all $i\in I_{00}$.
3 Solution Method
To solve (MPEC), traditional methods include relaxation method [16, 18, 19, 29, 37, 41], duality method [26, 25], exact penalization approach [38], stochastic method [34, 39]. We refer to [10] for review solving (MPEC). Keeping in mind that our aim is to design an efficient numerical algorithm to solve (MPEC), which is software-independent, the complementarity constraints are the first issue that we want to tackle. Smoothing technique has been widely used in solving many problems such as sparse optimization problems [2, 3] and nonsmooth optimization problems [28]. In this section, we solve (MPEC) by introducing smoothing functions to deal with the complementarity constraints in (MPEC). We will also investigate the relation of the smoothing problem and (MPEC).
In (MPEC), $0\leq H(v)\perp G(v)\geq 0$ are linear complementarity constraints. Therefore, from the optimization point of view, there are various methods to tackle complementarity constraints to their smoothing equivalent system to make the computation process more efficient [14, 30, 32, 33]. Different smoothing functions have been proposed and studied [4, 5, 11]. Here we choose the popular Fischer-Burmeister (FB) smoothing function [11], which is defined by
$\phi_{\epsilon}(a,b)=0$ if and only if $a>0,b>0,ab=\frac{\epsilon^{2}}{2}$. For $s\in\mathbb{R}^{m},\ t\in\mathbb{R}^{m}$, define
$\Phi_{\epsilon}(s,t)=\left[\phi_{\epsilon}(s_{1},t_{1}),\ \cdots,\ \phi_{\epsilon}(s_{m},t_{m})\right]^{\top}\in\mathbb{R}^{m}.$
The FB smoothing function has many nice properties useful in algorithmic development for semidefinite complementarity problems and nonsmooth equations involving the semidefinite conic complementarity condition. It is known from [42] that this function is strongly semismooth, which plays a fundamental role in the analysis of the quadratic convergence of Newton-type methods for solving such nonsmooth equation systems, see [36] for instance. Furthermore, it is proved by [40] that the gradient mapping of the squared norm of the FB smoothing function is Lipschitz continuous.
Applying $\Phi(\cdot,\cdot)$ to the complementarity constraints in (MPEC), we reach the smoothing version of (MPEC) as follows ($\epsilon>0$), which is a essentially a nonlinear programming problem (NLPϵ) with equality constraints:
$\begin{array}[]{cl}\min\limits_{v\in\mathbb{R}^{m+1}}&f(v)\\ \hbox{s.t.}&\Phi_{\epsilon}(G(v),H(v))=0.\end{array}$ (NLPϵ)
Here we would like to make a comparison between (MPEC) and (NLPϵ). In terms of variables, the two problems have the same scale of variables, which is $m+1$. However, the number of constraints in (NLPϵ) is only one third of that in (MPEC), which is much smaller. Moreover, the constraints in (NLPϵ) are equality constraints, which are much easier to deal with in designing algorithms.
[…]
Algorithm 1 Smoothing Method for (MPEC)
1:Choose $\epsilon_{0}\geq\epsilon_{\min}>0$, $\kappa\in(0,1)$, $t:=0$.
2:while $\epsilon_{t}\geq\epsilon_{\min}$, do
[…]
4: Let $\epsilon_{t+1}:=\kappa\epsilon_{t}$, and $t:=t+1$.
5:end while
6:Return the final iterate $v_{\rm opt}:=v^{t}$.
Here we would like to highlight the advantage of the proposed algorithm. Firstly, it is software independent, which do not rely on other packages. Additionally, by smoothing the complementarity constraints, the number of constraints is reduced to one third, which is much smaller than that solved by other state-of-art methods, and also contributes to the speed up of our algorithm.
The following theorem shows that under the fulfillment of MPEC-LICQ, any accumulation point of the smoothing algorithm is a C-stationary point of (MPEC).
Theorem 3.1.
[…]
Proof.
[…]
$\begin{array}[]{ll}\textbf{0}&=\nabla f(v^{t})-\sum_{i=1}^{m}\lambda^{(t)}_{i}\nabla\phi_{\epsilon_{t}}(G_{i}(v^{t}),H_{i}(v^{t}))\\ &=\nabla f(v^{t})-\sum_{i=1}^{m}\lambda^{(t)}_{i}\left(w^{G}_{i}(v^{t})\nabla G_{i}(v^{t})+w^{H}_{i}(v^{t})\nabla H_{i}(v^{t})\right),\end{array}$ (4)
where $i\in[m]$, and $\forall\epsilon_{t}>0$,
$\begin{array}[]{c}w^{G}_{i}(v^{t})=1-\dfrac{G_{i}(v^{t})}{\sqrt{G^{2}_{i}(v^{t})+H_{i}^{2}(v^{t})+(\epsilon_{t})^{2}}}\in(0,1],\\ w^{H}_{i}(v^{t})=1-\dfrac{H(v^{t})}{\sqrt{G^{2}_{i}(v^{t})+H_{i}^{2}(v^{t})+(\epsilon_{t})^{2}}}\in(0,1].\end{array}$
Together with $\{\epsilon_{t}\}\searrow 0$, and that $v^{*}$ is an accumulation point of $v^{t}$, by taking subsequence if necessary, assume that $w^{G}_{i}(v^{t})\to a_{i}$, $w^{H}_{i}(v^{t})\to b_{i}$. By defining
$\begin{array}[]{c}I_{00}^{1G}=\{i\in I_{00}\ |\ a_{i}=0\},\ I_{00}^{2G}=\{i\in I_{00}\ |\ a_{i}=1\},\ I_{00}^{3G}=\{i\in I_{00}\ |\ a_{i}\in(0,1)\},\\ I_{00}^{1H}=\{i\in I_{00}\ |\ b_{i}=0\},\ I_{00}^{2H}=\{i\in I_{00}\ |\ b_{i}=1\},\ I_{00}^{3H}=\{i\in I_{00}\ |\ b_{i}\in(0,1)\}.\end{array}$
it holds that
$\begin{array}[]{cc}a_{i}\in\begin{cases}\{0\},\ i\in I_{+0},\\ \{1\},\ i\in I_{0+},\\ \{0\},\ i\in I_{00}^{1G},\\ \{1\},\ i\in I_{00}^{2G},\\ (0,1),\ i\in I_{00}^{3G},\end{cases}&b_{i}\in\begin{cases}\{1\},\ i\in I_{+0},\\ \{0\},\ i\in I_{0+},\\ \{0\},\ i\in I_{00}^{1H},\\ \{1\},\ i\in I_{00}^{2H},\\ (0,1),\ i\in I_{00}^{3H}.\end{cases}\end{array}$
Assume for contradiction that the sequence $\left\{\lambda^{(t)}\right\}$ is unbounded, then one can find a subsequence $K$ such that the $\frac{\lambda^{(t)}}{\|\lambda^{(t)}\|}$ converges to some vector, denoted as $\hat{\lambda}\neq 0$.
Denote $\hat{I}_{+}:=\{i\mid\hat{\lambda}_{i}>0\}$, $\hat{I}_{-}:=\{i\mid\hat{\lambda}_{i}<0\}$, $\hat{I}_{0}:=\{i\mid\hat{\lambda}_{i}=0\}$. By taking subsequence if necessary, it holds that
$\dfrac{\lambda^{(t)}}{\|\lambda^{(t)}\|}w^{G}_{i}(v^{t})\to\hat{\gamma_{i}}=\hat{\lambda}_{i}a_{i},\ \dfrac{\lambda^{(t)}}{\|\lambda^{(t)}\|}w^{H}_{i}(v^{t})\to\hat{\nu_{i}}=\hat{\lambda}_{i}b_{i}.$